Please solve it by hand or typing, don't use EXCEL.

1. XYZ stock price and dividends history are as follows Year Beginning-of-year price Dividend paid at year- $100 $110 $90 $95 end $4 $4 $4 $4 2005 2006 2007 2008 An investor buys three shares of XYZ at the beginning of 2005, buys another two shares at the beginning of 2006, sells one share at the beginning of 2007, and sells all four remaining shares at the beginning of 2008. Hint: When the investor buys three shares at the beginning of 2005, she cannot receive dividends paid at the end of 2004. She can receive dividend payments on these three shares at the end of 2005 A. Computer HPR for each period (05-06, 06-07, and 07-08) (6pts) B. Compute the arithmetic return. (3pts) C. Compute the geometric return. (3pts) D. Computer the dollar-weighted rate of return. Draw the time line to show cash flows. Write down the formula and calculate the dollar-weighted rate of return. The time line includes four dates corresponding to the turns of the year for January 1, 2005 to January 1, 2008. (8pts)
Answer #1

a) HPR

A B C D=B-A E=(D+C)/A
Year Beigining Price Ending Price Dividend / Share Capital Gain Return = (Capital Gain + Dividend)/Price
2005-06 100 110 4 10 14.00% (110-100+4)/100
2006-07 110 90 4 -20 -14.55% (90-110+4)/110
2007-08 90 95 4 5 10.00% (95-90+4)/90

b) Arithmetic  return:

AR = (HPR1 + HPR2 + HPR3)/ 3 = (14% - 14.55% + 10%) / 3 = 3.15%

c) Geometric Return

GR ={(1+HPR1)*(1+HPR2)*(1+HPR3)} ^(1/3) - 1

GR = (1+0.14)*(1-0.1455)*(1+0.1)-1

GR= 2.33  %

d)

Cash Flow T CF Calculation Description
01-01-2005 0 -300 (-100*3) Purchase of 3 shares
01-01-2006 1 -208 (-110*2+4*3) Purchase of 2 shares plus dividend on 3 shares held
01-01-2007 2 110 (90*1+4*5) Dividend on 5 Shares plus sale of 1 share @ 90
01-01-2008 3 396 (95*4+4*4) Dividend on 4 Shares plus sale of 1 share @ 95

Dollar-time weighted return

0 = -300 + -208 / (1+r) + 110 / (1+r)2 + 396/(1+r)3

r= - 0.17%